OpenStudy (anonymous):

Matthew is training to participate in a half marathon. The table below shows the number of miles he runs in weeks 1 and 2. Week Number of Miles Run 1 12.5 2 13.5 Each week, Matthew increases the number of miles he runs in the previous week by the same percent. How many miles does Matthew run in week 3? Do not round your answer.

OpenStudy (triciaal):

one approach change in miles between wk 2 and wk 1 = 13.5 - 12.5 =1 change as a percentage = 1/12.5 * 100 =g % (remember % means per 100) change between wk 2 and wk 3 is the same percentage wk 2 = 13.5 g% of 13.5 = ? wk 3 = wk 2 + increase